# Optimization via Quantum Preconditioning

### v1

@@ III.1 Random $3$-Regular Graph Maximum-Cut Problems
III.1.1 Classical Simulated Annealing Solver Baseline
We investigate the average performance of SA [6] on hundreds of random $N$-variable $3$-regular graph maximum-cut problems. SA is a physics-inspired solver treating an objective function such as Eq. (1) as a classical Ising model from which one wants to find the ground state. The search is based on a Metropolis-Hastings Markov chain Monte-Carlo algorithm with a Boltzmann distribution [50, 51]. It starts at a high temperature reduced throughout the iterations according to a predefined schedule such that one samples low-energy states, and ideally the ground state (see Appendix D). The objective function of the maximum-cut problem is closely related to that of Eq. (1) as it seeks to maximize the cut number
@@ III.1.1 Classical Simulated Annealing Solver Baseline
where $W_{ij}=1$ if there is an edge between vertices $i$ and $j$ and zero otherwise. We evaluate in Fig. 1c the average approximation ratio $\alpha$ as a function of the solver’s run-time for $N=4,096$ variables. We estimate and report the run-time to get to $\alpha=99.9\%$ as a function of $N$ in Fig. 1b.
@@ III.1.2 Performing Quantum Preconditioning
Having defined a classical baseline, we now quantum-precondition the problems using exact state vector emulations. We employ a light-cone decomposition technique [45] to access two-point correlations $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}$ of Eq. (2) on problems much lager than qubits naïvely accessible.
A $p$-layer QAOA circuit induces quantum coherence around each qubit and its $p$-nearest neighbors. Therefore, one can estimate a two-point correlation between two qubit $i$ and $j$ via the QAOA by only considering the subset of qubits resulting from the intersection of their respective light cones. If the lights cone do not intersect, then $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}=\langle\hat{Z}_{i}\rangle_{p}\langle\hat{Z}_{j}\rangle_{p}=0$ due to the global sign-flip symmetry of Eq. (5). The topology of $3$-regular graphs is such that intersecting light cones have at most $1+6(2^{p}-1)$ qubits, making $p\leq 2$ readily accessible through standard circuit emulation methods.
@@
Figure 2: Average performance for the maximum-cut problem on $N$-variable random $3$-regular graphs via the classical BM solver based on the original and quantum-preconditioned problems. (A) Average run-time for BM to get to an approximation ratio of $99.9\%$ as a function of $N$. (B) Average approximation ratio versus BM run-time for $N=4,096$. Each data point is averaged over $200$ randomly generated problem instances. Run-times corresponds to a 64Gb MacBook Pro with an Apple M1 Max chip. Error bars indicate the standard error of the mean.
Quantum preconditioning increases the number of nonzero terms $n$ in the problem because correlations extend beyond nearest-neighbor: There are $n=1.5N$ terms in the original $3$-regular graph problem, $n\simeq 4.5N$ at $p=1$, and $n\simeq 22.5N$ at $p=2$ in the large $N$ limit (see Appendix E). Although classical solvers typically have a run-time proportional to the number of terms $n$ (see Appendix H), we show in the following that the increase in the number of terms is not a bottleneck for quantum preconditioning to deliver an advantage.
@@ III.1.3 Quantum-Preconditioned Simulated Annealing Solver
We report in Fig. 1 the performance of quantum preconditioning on SA for hundreds of random $3$-regular graph maximum-cut problems. We find that SA gets to an approximation ratio of $99.9\%$ about one order of magnitude faster when working with the quantum-preconditioned input rather than the original problem. For the largest problem sizes considered, we also observe a separation between $p=1$ and $p=2$, where increasing the circuit depth further accelerates the convergence of the classical solver. At $p\to+\infty$, we anticipate SA to trivially converge to the optimal solution.
At a fixed number of variables ($N=4,096$), Fig. 1c shows that quantum preconditioning can deliver a benefit across a range of approximation ratio and run-time targets. As the run-time increases, we observe that SA on the $p=2$ quantum preconditioned problem saturates to an approximation ratio $\alpha<1$. This is because as mentioned earlier, at finite $p$, the optimal solutions of the original and quantum preconditioned problems are generically different. However, it is interesting to see that in practice, their respective optimal solutions are spectrally close with the solver attaining $\alpha\simeq 99.96\%$ at $p=2$. We provide additional data for other values of $N$ in Appendix G.
@@ III.1.4 Quantum-Preconditioned Burer-Monteiro Solver
Next, we consider quantum preconditioning in the context of the BM solver [8, 9]. The BM solver (see Appendix D) is inspired by semidefinite programming methods: It relaxes the discrete nature of the variables in the objective function, solves this modified problem, and rounds the solution back to the valid domain at the end. In practice, the BM algorithm is one of the best heuristics to date for the maximum-cut problem [12]. Depending on the problem size and the performance target, it may be a better choice than SA.
@@
The performance boost offered by quantum preconditioning and observed on solvers as different as SA and BM suggests the wide applicability of the technique. Especially, the quantum relax-and-round solver [49, 45] is another example of quantum preconditioning via a classical relax-and-round solver inspired by semidefinite programming methods. In that case, one can prove that the performance at $p=1$ matches that of the classical counterpart and increases with $p$ on classes of problems, including Sherrington-Kirkpatrick spin glasses and random $3$-regular graph maximum-cut problems [49, 45].
@@ III.2.1 Problem and Quantum Preconditioning
We now turn our attention to another class of problems: Sherrington-Kirkpatrick (SK) spin glasses [43]. The goal is to minimize an objective function of the form of Eq. (1), where the adjacency matrix $\mathsf{W}$ encoding the problem is drawn from the Gaussian orthogonal ensemble. Namely, its entries are independent and identically distributed random variables from a normal distribution of zero mean and unit width $\mathsf{W}_{ij}\sim\mathcal{N}(0,1)$ with $\mathsf{W}_{ij}=\mathsf{W}_{ji}$. SK problem instances are quite different from the random $3$-regular graph maximum-cut problems considered previously and constitute another standard optimization benchmark [53, 27, 54, 55, 56]. First, they are dense with $O(N^{2})$ terms instead of $O(N)$ for $N$-variable problems. Second, they are weighted problems with $\mathsf{W}_{ij}\neq 1$.
We perform quantum preconditioning at $p=1$, where one can compute semi-analytically the correlation matrix of Eq. (2) via a back-propagation technique [49] (reproduced in Appendix F for completeness). We use optimal QAOA angles $\gamma=1/2\sqrt{N}$ and $\beta=\pi/8$ from the $N\to+\infty$ limit [27], thus skipping variational optimization. Unlike the $3$-regular graph maximum-cut problems, the resulting quantum-preconditioned SK spin glasses conserve the number of terms and the same statistics as the original $\mathsf{W}_{ij}$ [49]. Emulating the QAOA at $p>1$ for large $N$ is currently not feasible.
@@ III.2.2 Performance
We show in Fig. 3 the performance of SA and the BM solver on $200$ randomly generated SK instances with $N=2,048$ variables (additional data at other values of $N$ is available in Appendix G). In particular, we plot the approximation ratio as a function of the number of iterations (Fig. 3a) and run-time (Fig. 3b) with the relation between the two discussed in Appendix H.
Figure 3: Average performance for $N=2,048$ Sherrington-Kirkpatrick spin glass problems via the classical simulated annealing (SA) and Burer-Monteiro (BM) solvers based on the original and quantum-preconditioned $(p=1)$ problems. (A) Average approximation ratio as a function of the number of iterations. (B) Average approximation ratio as a function of the run-time. Each data point is averaged over $200$ randomly generated problem instances. Run-times corresponds to a 64Gb MacBook Pro with an Apple M1 Max chip. Error bars indicate the standard error of the mean.
Figure 4: Performance of quantum preconditioning for computing the maximum power exchange section of an energy grid optimization problem. (A) Representation of the grid energy optimization problem considered from the state of South Carolina in the United States of America. Vertices ($N=180$) and edges ($n=226$) correspond to buses and lines, respectively. (B) Average approximation ratio as a function of the number of iterations via simulated annealing (SA). (C) Average approximation ratio as a function of run-time via SA. (D) Average approximation ratio as a function of the number of iterations via the Burer-Monteiro (BM) solver. (E) Average approximation ratio as a function of run-time via the BM solver. Run-times are based on a 64Gb MacBook Pro with an Apple M1 Max chip. Each data point is averaged over $10^{4}$ samples. Error bars indicate the standard error of the mean.
We observe that for a large iteration number (or run-time), the approximation ratio obtained via the quantum-preconditioned problem saturates to $\alpha\simeq 99.3\%$. This, again, indicates that it does not share the same optimal solution as the original problem. At a small iteration number, we find that at a fixed number of iterations, the classical solver benefits from working with the preconditioned problem as it achieves a higher approximation ratio. This leads to a crossing of performance and an advantage window for quantum preconditioning on SK problems for both SA and the BM solvers.
The overhead of about two seconds for the SA data in Fig. 3b originates from the $O(N^{2})$ heuristic used for setting the temperature schedule (see Appendix D). It is absent in the case of the BM solver which doesn’t have any tuned hyperparameters. The crossing of performance happens at slightly different approximation ratio values for SA and the BM algorithm. This can be understood as both solvers returning different distribution of solutions.
III.3 Power Grid Optimization Problem: The Maximum Power Exchange Section
@@ III.3.1 Problem Definition
The last example considered is a real-world power grid optimization problem seeking to compute the maximum power exchange section (MPES) of an energy network [57, 58, 59, 60]. The MPES is a metric that informs on the health and the power delivery capability of the energy grid. We consider the realistic $500$-bus power system dataset “ACTIVSg500” mimicking the energy grid of the state of South Carolina in the United States of America [44], see Fig. 4a. The dataset was designed as part of the ARPA-E’s GRID DATA program to be similar to the actual electric grid while containing no confidential critical energy infrastructure information [61, 62, 63, 64, 65].
@@
with $\mathsf{W}_{ji}=\mathsf{W}_{ij}$ and $\mathsf{W}_{ij}=0$ if buses $i$ and $j$ are not connected. $\mathsf{R}_{ij}$ and $\mathsf{X}_{ij}$ correspond to the line resistance and reactance, respectively. Their values are provided as part of the dataset.
The original problem contains $N=410$ variables and $n=463$ terms. The corresponding graph problem contains dangling tree branches, which we remove. Because their topology is nonfrustrating, it is trivial to find the optimal value for the variables on those branches once a solution for the rest of the problem is found: One simply sets the value of a variable $z_{j}=\pm 1$ such that $\mathsf{W}_{ij}z_{i}z_{j}$ is locally minimized. The remaining problem has $N=201$ variables and $n=254$ terms. However, one finds that it actually consists of two disconnected and independent subgraphs. The first one contains $N=180$ nodes and $n=226$ edges. The second contains $N=21$ nodes and $n=28$ edges and can, therefore, be solved via brute force enumeration. In the following, we solve the largest of the two subproblems and trivially reconstruct a solution $\boldsymbol{z}$ to the original problem of Eq. (6).
The optimal contribution of the dangling tree branches to the global objective value is $\simeq 50,778.093880$, the optimal contribution of the $N=21$ nodes problem to the global objective value is $\simeq 3,728.413221$, and we estimate that the global optimal solution has an objective value $C_{\textrm{MPES}}(\boldsymbol{z}_{\textbf{opt}})\simeq 61,606.078827$.
@@ III.3.2 Quantum Preconditioning
The $N=180$ graph problem is sufficiently sparse that we can employ the light cone technique introduced in the context of the random $3$-regular graph maximum-cut problems in Sec. III.2.1. The graph has an average and maximum degree of $\simeq 2.51$ and $7$, respectively. At $p=1$ and $p=2$, the light-cone induced subgraphs for building the correlation matrix have at most $14$ and $27$ vertices, respectively. Both numbers are within the ability of classical state vector emulators, which we use for preconditioning.
Unlike the previous graph problems investigated, near-optimal QAOA angles are not tabulated for this problem. As such, we perform a variational search of the QAOA angles extremizing the expectation value of the objective function $\langle\hat{C}\rangle_{p}$. We use the Broyden-Fletcher-Goldfarb-Shanno (BFGS) [66, 67, 68, 69] algorithm starting from random initial angles a few hundred times and consider the best output for computing the correlation matrix. We report the QAOA angles in Appendix A. The original and preconditioned problems have $n=226$, $n=600$ $(p=1)$, and $n=1,625$ $(p=2)$ terms, respectively.
@@ III.3.3 Performance
We display the average approximation as a function of the number of iterations and run-time for SA and the BM solvers in Fig. 4. We compare the performance of these solvers on the original and the preconditioned problems. We evaluate the average approximation ratio by sampling $10^{4}$ solutions from the classical solvers for each iteration number. We observe that for this problem in the range considered, the BM algorithm is more efficient than SA as it can reach a much higher approximation ratio for the same fixed run-time.
For a given number of iterations, we find regions where the solvers on the preconditioned problems yield a higher approximation ratio than with the original one. This defines windows for an advantage. At a small number of iterations, we observe that the window is enlarged from $p=1$ to $p=2$. Such a gain from increasing the amount of preconditioning was also observed for the random $3$-regular graph maximum-cut problems. When substituting the number of iterations for the run-time on a 64Gb MacBook Pro with an Apple M1 Max chip, there still exists an advantage window despite the quantum preconditioned problems having more terms than the original one (see Appendix A). Finally, we note that $p=1,2$ preconditioned problems can access approximation ratios $\alpha\gtrsim 99.99\%$ and that they do not share the same optimal solutions as the original problem as exemplified by the saturation in approximation ratio observed in Fig. 4d. In addition to classical state vector emulations, we precondition the problem using an experimental implementation of the QAOA at $p=1$, and solve the preconditioned problem. Results of this experiment are shown in App. C.
The existence of an advantage window on a real-world problem highlights the potential wide applicability of the proposed quantum preconditioning method.
IV Quantum Preconditioning Budget and Prospects for a Quantum Advantage
We have shown that quantum preconditioning problems can benefit state-of-the-art classical solvers. While this is a nontrivial result on its own, a practical advantage can only be claimed if one also accounts for the preconditioning time itself—see, e.g., Fig. 1b defining a preconditioning time budget for random $3$-regular graph maximum-cut problems via SA.

### v2

@@ III.1 Random $3$-regular graph maximum-cut problems
III.1.1 Classical simulated-annealing-solver baseline
We investigate the average performance of SA [6] on hundreds of random $N$-variable 33-regular graph maximum-cut problems. SA is a physics-inspired solver treating an objective function such as Eq. (1) as a classical Ising model from which one wants to find the ground state. The search is based on a Metropolis-Hastings Markov chain Monte Carlo algorithm with a Boltzmann distribution [49, 50]. It starts at a high temperature, which is reduced throughout the iterations according to a predefined schedule such that one samples low-energy states, and ideally the ground state (see Appendix E). The objective function of the maximum-cut problem is closely related to that of Eq. (1), as it seeks to maximize the cut number
@@ III.1.1 Classical simulated-annealing-solver baseline
where $W_{ij}=1$ if there is an edge between vertices $i$ and $j$ and zero otherwise. In Fig. 1c, we evaluate the average approximation ratio $\alpha$ as a function of the solver’s run-time for $N=4,096$ variables. We estimate and report the run-time to get to $\alpha=99.9\%$ as a function of $N$ in Fig. 1b.
@@ III.1.2 Performing quantum preconditioning
Having defined a classical baseline, we now quantum precondition the problems using exact state-vector emulations. We employ a light-cone decomposition technique [44] to access two-point correlations $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}$ of Eq. (2) on problems much larger than qubits naïvely accessible.
A $p$-layer QAOA circuit induces quantum coherence around each qubit and its $p$-nearest neighbors. Therefore, one can estimate a two-point correlation between two qubits $i$ and $j$ via the QAOA by only considering the subset of qubits resulting from the intersection of their respective light cones. If the light cones do not intersect, then $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle_{p}=\langle\hat{Z}_{i}\rangle_{p}\langle\hat{Z}_{j}\rangle_{p}=0$ due to the global sign-flip symmetry of Eq. (5). The topology of $3$-regular graphs is such that intersecting light cones have at most $1+6(2^{p}-1)$ qubits, making $p\leq 2$ readily accessible through standard circuit-emulation methods.
@@
Figure 2: Average performance for the maximum-cut problem on $N$-variable random $3$-regular graphs via the classical BM solver based on the original and quantum-preconditioned problems. (A) Average run-time for BM to get to an approximation ratio of $99.9\%$ as a function of $N$. (B) Average approximation ratio versus BM run-time for $N=4,096$. Each data point is averaged over $200$ randomly generated problem instances. Run-times correspond to a 64GB MacBook Pro with an Apple M1 Max chip. Error bars indicate the standard error of the mean.
Quantum preconditioning increases the number of nonzero terms $n$ in the problem because correlations extend beyond nearest-neighbor: There are $n=1.5N$ terms in the original $3$-regular graph problem, $n\simeq 4.5N$ at $p=1$, and $n\simeq 22.5N$ at $p=2$ in the large-$N$ limit (see Appendix G). Although classical solvers typically have a run-time proportional to the number of terms $n$ (see Appendix K), we show in the following that the increase in the number of terms is not a bottleneck for quantum preconditioning to deliver an advantage.
@@ III.1.3 Quantum-preconditioned simulated annealing solver
In Fig. 1, we report the performance of quantum preconditioning on SA for hundreds of random $3$-regular graph maximum-cut problems. We find that SA gets to an approximation ratio of $99.9\%$ about one order of magnitude faster when working with the quantum-preconditioned input rather than the original problem. For the largest problem sizes considered, we also observe a separation between $p=1$ and $p=2$, where increasing the circuit depth further accelerates the convergence of the classical solver. At $p\to+\infty$, we anticipate that SA will trivially converge to the optimal solution.
At a fixed number of variables ($N=4,096$), Fig. 1c shows that quantum preconditioning can deliver a benefit across a range of approximation-ratio and run-time targets. As the run-time increases, we observe that SA on the $p=2$ quantum preconditioned-problem saturates to an approximation ratio $\alpha<1$. This is because, as mentioned earlier, at finite $p$, the optimal solutions of the original and quantum-preconditioned problems are generically different; however, it is interesting to see that in practice, their respective optimal solutions are spectrally close with the solver attaining $\alpha\simeq 99.96\%$ at $p=2$. We provide additional data for other values of $N$ in Appendix J.
@@ III.1.4 Quantum-preconditioned Burer-Monteiro solver
Next, we consider quantum preconditioning in the context of the BM solver [8, 9]. The BM solver (see Appendix E) is inspired by semidefinite programming methods: it relaxes the discrete nature of the variables in the objective function, solves this modified problem, and rounds the solution back to the valid domain at the end. In practice, the BM algorithm is one of the best heuristics to date for the maximum-cut problem [12]. Depending on the problem size and the performance target, it may be a better choice than SA.
@@
The performance boost offered by quantum preconditioning and observed on solvers as different as SA and BM suggests the wide applicability of the technique. In particular, the quantum relax-and-round solver [48, 44] is another example of quantum preconditioning via a classical relax-and-round solver inspired by semidefinite programming methods. In that case, one can prove that the performance at $p=1$ matches that of the classical counterpart and increases with $p$ on classes of problems—including Sherrington-Kirkpatrick spin glasses and random $3$-regular graph maximum-cut problems [48, 44].
@@ III.2.1 Problem and quantum preconditioning
We now turn our attention to another class of problems: Sherrington-Kirkpatrick (SK) spin glasses [42]. The goal is to minimize an objective function of the form of Eq. (1), where the adjacency matrix $\mathsf{W}$ encoding the problem is drawn from the Gaussian orthogonal ensemble. Namely, its entries are independent and identically distributed random variables from a normal distribution of zero mean and unit width $\mathsf{W}_{ij}\sim\mathcal{N}(0,1)$, with $\mathsf{W}_{ij}=\mathsf{W}_{ji}$. SK problem instances are quite different from the random $3$-regular graph maximum-cut problems considered previously and constitute another standard optimization benchmark [52, 27, 53, 54, 55]. First, they are dense, with $O(N^{2})$ terms instead of $O(N)$ for $N$-variable problems. Second, they are weighted problems with $\mathsf{W}_{ij}\neq 1$.
We perform quantum preconditioning at $p=1$, where one can compute semianalytically the correlation matrix of Eq. (2) via a back-propagation technique [48] (reproduced in Appendix H for completeness). We use optimal QAOA angles $\gamma=1/2\sqrt{N}$ and $\beta=\pi/8$ from the $N\to+\infty$ limit [27], thus skipping variational optimization. Unlike the $3$-regular graph maximum-cut problems, the resulting quantum-preconditioned SK spin glasses conserve the number of terms and the same statistics as the original $\mathsf{W}_{ij}$ [48]. Emulating the QAOA at $p>1$ for large $N$ is currently not feasible.
@@ III.2.2 Performance
In Fig. 3, we show the performance of SA and the BM solver on $200$ randomly generated SK instances with $N=2,048$ variables (additional data at other values of $N$ are available in Appendix J). In particular, we plot the approximation ratio as a function of the number of iterations (Fig. 3a) and run-time (Fig. 3b) with the relation between the two discussed in Appendix K.
Figure 3: Average performance for $N=2,048$ Sherrington-Kirkpatrick spin-glass problems via the SA and Burer-Monteiro (BM) solvers based on the original and quantum-preconditioned $(p=1)$ problems. (A) Average approximation ratio as a function of the number of iterations. (B) Average approximation ratio as a function of the run-time. Each data point is averaged over $200$ randomly generated problem instances. Run-times correspond to a 64GB MacBook Pro with an Apple M1 Max chip. Error bars indicate the standard error of the mean.
Figure 4: Performance of quantum preconditioning for computing the maximum power exchange section of an energy-grid optimization problem. (A) Representation of the energy-grid optimization problem considered from the state of South Carolina in the United States of America. Vertices ($N=180$) and edges ($n=226$) correspond to buses and lines, respectively. (B) Average approximation ratio as a function of the number of iterations via SA. (C) Average approximation ratio as a function of run-time via SA. (D) Average approximation ratio as a function of the number of iterations via the BM solver. (E) Average approximation ratio as a function of run-time via the BM solver. Run-times are based on a 64GB MacBook Pro with an Apple M1 Max chip. Each data point is averaged over $10^{4}$ samples. Error bars indicate the standard error of the mean.
We observe that for a large number of iterations (or run-time), the approximation ratio obtained via the quantum-preconditioned problem saturates to $\alpha\simeq 99.3\%$. This, again, indicates that it does not share the same optimal solution as the original problem. For a small number of iterations, we find that the classical solver benefits from working with the preconditioned problem, achieving a higher approximation ratio at the same iteration count. This leads to a crossing of performance and an advantage window for quantum preconditioning on SK problems for both SA and BM solvers.
The overhead of about two seconds for the SA data in Fig. 3b originates from the $O(N^{2})$ heuristic used for setting the temperature schedule (see Appendix E). It is absent in the case of the BM solver, which does not have any tuned hyperparameters. The crossing of performance happens at slightly different approximation ratio values for SA and the BM algorithm. This can be understood as both solvers returning different distributions of solutions.
III.3 Power-grid optimization problem: The maximum power exchange section
@@ III.3.1 Problem definition
The last example considered is a real-world power-grid optimization problem seeking to compute the maximum power exchange section (MPES) of an energy network [56, 57, 58, 59]. The MPES is a metric that provides information about the health and the power-delivery capability of the energy grid. We consider the realistic $500$-bus power system dataset ACTIVSg500, which mimics the energy grid of the state of South Carolina in the United States of America [43]; see Fig. 4a. The dataset was designed as part of the ARPA-E’s GRID DATA program to be similar to the actual electric grid while containing no confidential information about critical energy infrastructure [60, 61, 62, 63, 64].
@@
with $\mathsf{W}_{ji}=\mathsf{W}_{ij}$ and $\mathsf{W}_{ij}=0$ if buses $i$ and $j$ are not connected. Further, $\mathsf{R}_{ij}$ and $\mathsf{X}_{ij}$ correspond to the line resistance and reactance, respectively. Their values are provided as part of the dataset.
The original problem contains $N=410$ variables and $n=463$ terms. The corresponding graph problem contains dangling tree branches, which we remove. Because their topology is nonfrustrating, it is trivial to find the optimal value for the variables on those branches once a solution for the rest of the problem is found: one simply sets the value of a variable $z_{j}=\pm 1$ such that $\mathsf{W}_{ij}z_{i}z_{j}$ is locally minimized. The remaining problem has $N=201$ variables and $n=254$ terms; however, one finds that it actually consists of two disconnected and independent subgraphs. The first contains $N=180$ nodes and $n=226$ edges; the second contains $N=21$ nodes and $n=28$ edges and can, therefore, be solved via brute-force enumeration. In the following, we solve the larger of the two subproblems and trivially reconstruct a solution $\boldsymbol{z}$ to the original problem of Eq. (6).
The optimal contribution of the dangling tree branches to the global objective value is approximately $50,778.093880$, the optimal contribution of the $N=21$ nodes problem to the global objective value is approximately $3,728.413221$, and we estimate that the global optimal solution has an objective value $C_{\textrm{MPES}}(\boldsymbol{z}_{\textbf{opt}})\simeq 61,606.078827$.
@@ III.3.2 Quantum preconditioning
The $N=180$ graph problem is sufficiently sparse that we can employ the light-cone technique introduced in the context of the random $3$-regular graph maximum-cut problems in Sec. III.1.2. The graph has an average and maximum degree of approximately $2.51$ and $7$, respectively. At $p=1$ and $p=2$, the light-cone-induced subgraphs for building the correlation matrix have at most $14$ and $27$ vertices, respectively. Both numbers are within the ability of classical state-vector emulators, which we use for preconditioning.
Unlike the previous graph problems investigated, near-optimal QAOA angles are not tabulated for this problem. As such, we perform a variational search for the QAOA angles extremizing the expectation value of the objective function $\langle\hat{C}\rangle_{p}$. We use the Broyden-Fletcher-Goldfarb-Shanno (BFGS) [65, 66, 67, 68] algorithm, starting from random initial angles a few hundred times and consider the best output for computing the correlation matrix. We report the QAOA angles in Appendix A. The original and preconditioned problems have $n=226$, $n=600$ $(p=1)$, and $n=1,625$ $(p=2)$ terms, respectively.
@@ III.3.3 Performance
We display the average approximation as a function of the number of iterations and run-time for SA and the BM solver in Fig. 4. We compare the performance of these solvers on the original and preconditioned problems. We evaluate the average approximation ratio by sampling $10^{4}$ solutions from the classical solvers for each number of iterations. We observe that for this problem in the range considered, the BM algorithm is more efficient than SA as it can reach a much higher approximation ratio for the same fixed run-time.
For a given number of iterations, we find regions where the solvers on the preconditioned problems yield a higher approximation ratio than with the original one. This defines windows for an advantage. At a small number of iterations, we observe that the window is enlarged from $p=1$ to $p=2$. Such a gain from increasing the amount of preconditioning was also observed for the random $3$-regular graph maximum-cut problems. When substituting the number of iterations for the run-time on a 64GB MacBook Pro with an Apple M1 Max chip, there is still an advantage window, despite the quantum-preconditioned problems having more terms than the original one (see Appendix G). Finally, we note that $p=1,2$ preconditioned problems can access approximation ratios $\alpha\gtrsim 99.99\%$ and that they do not share the same optimal solutions as the original problem as exemplified by the saturation in approximation ratio observed in Fig. 4d. In addition to classical state-vector emulations, we precondition the problem using an experimental implementation of the QAOA at $p=1$ and solve the preconditioned problem. Results of this experiment are shown in App. D.
The existence of an advantage window for a real-world problem highlights the potential wide applicability of the proposed quantum-preconditioning method.
IV Quantum-Preconditioning Budget and Prospects for a Quantum Advantage
We have shown that quantum-preconditioning problems can benefit state-of-the-art classical solvers. While this is a nontrivial result on its own, a practical advantage can only be claimed if one also accounts for the preconditioning time itself—see, e.g., Fig. 1b defining a preconditioning time budget for random $3$-regular graph maximum-cut problems via SA.